OpenStudy (anonymous):

assume that 30% of all plants in a field are infested with aphids. suppose that you pick 2 plants at random. what is the probability that exactly one of them carried aphids?

OpenStudy (anonymous):

It depends on how many plants are in the field

OpenStudy (anonymous):

that's the question. i got 0.49

OpenStudy (anonymous):

0.3= infested and 0.7=not infested

OpenStudy (anonymous):

so i did 1-0.7^2=1-0.51=0.49

OpenStudy (anonymous):

So how many plants are there in total in the field. This is needed to calculate the probability

OpenStudy (anonymous):

it doesn't say how many plants are in the field

OpenStudy (anonymous):

it gives u 30% of the plants are infested which means that 70% of the plants are not infested

OpenStudy (anonymous):

so in that field which an unknown amount of plants it's asking for the probability that exactly one of them carried aphids.

OpenStudy (anonymous):

Is this out of a text book?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it's an even number question and has no answers so i just wanted to be sure that i'm doing it right

OpenStudy (anonymous):

I'm writing a program to test different plant numbers and get the probability...

OpenStudy (anonymous):

ok. thank you :)

OpenStudy (anonymous):

Definitely needs plant numbers: In a field with 10 plants the probability is 0.4666666666666667 with 21/45 combinations carrying 1 aphid infested plant In a field with 100 plants the probability is 0.42424242424242425 with 2100/4950 combinations carrying 1 aphid infested plant In a field with 100 plants the probability is 0.42042042042042044 with 210000/499500 combinations carrying 1 aphid infested plant So it depends. Would you like to verify the code (if you understand it)

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Sure you would like to verify it? OK: public class PlantsInField { public static void main(String[] args) { int limit = 1000; double prob = 0; double total = calcTotal(limit); for (int A = 1; A <= limit; A++) { for (int B = A + 1; B <= limit; B++) { int x = 0; int y = 0; if (A <= limit * 0.3) x = 1; if (B <= limit * 0.3) y = 1; if ((x^y) == 1) { prob += 1; } } } System.out.println(prob+"/"+total); prob /= total; System.out.println(prob); } static int calcTotal(int N) { int n = N; int product = 1; for (; n >= N - 1; n--) { product *= n; } return product / 2; } }

OpenStudy (anonymous):

that's fine. i don't understand programming. Thank you though :)

OpenStudy (anonymous):

Sure thing